Scenario 4

A small college has historically found that 17% of its students withdraw without completing the introductory statistics course. Twenty-two students have enrolled in the course this semester. Use Excel for this scenario. (Round probabilities to four decimals)

1) What is the probability that 5 or fewer will withdraw.

2) What is the probability that exactly 3 withdraw?

3) What is the probability that 6 or more withdraw?

4) How many students are expected to withdraw? (Two decimals)
Answer #1

Sceario 4

Here sample size = n = 22

sample proportion = p = 0.17

(1) P(x 5) = BINOMDIST(5; 22 ; 0.17; TRUE) = 0.8424

(2) P(x = 3) = BINOMDIST(3; 22 ; 0.17; False) = 0.2195

(3) P(x 6) = 1 - P(x 5) = 1- 0.8424 = 0.1576

(4) E[X] = np = 22 * 0.17 = 3.74